OpenStudy (anonymous):

Look at the formula for the volume of a cylinder. V = pi r2h Which is a solution for h?

OpenStudy (anonymous):

h=V/pi r2

OpenStudy (anonymous):

h=V/(pi r^2)

OpenStudy (anonymous):

h=v/2r^2pi

OpenStudy (anonymous):

h=V/[pi r^2]

OpenStudy (anonymous):

\[h=V/\pi r^2\]